Please help solve for 'x' 4x^2=5x-7<12x^2+16x-10
Answer by Pumuky(2)

4x^2=5x-7 < 12x^2+16x-10
4x^2-5x+7=0 < 12x^2+16x-10
I dont understand your question. The first equation have a solution for x but the second equation is a funtcion. Two point are not minor o higher that one function. They are two different things.
If you need the area between two function you can contniue with:
4x^2-5x+7=0 < 12x^2+16x-10
4x^2-5x+7 < 12x^2+16x-10
0 =< 12x^2+16x-10-4x^2+5x-7
0 =< 8x^2+21x-17
The roots are x=-3.27 and x=0.64 and the function are positive if x is less that -3.27 or higher that 0.64
Espero que te sirva. I hope that it will be useful